# Non-autonomous iteration of polynomials in the complex plane

### v2

@@ 1. Introduction
Behavior of iterates $p^{n}:=p^{\circ n}$ of a polynomial $p:\mathbb{C}\longrightarrow\mathbb{C}$ has been intensely studied since the nineteenth century. For an introduction to the subject see e.g. [7]. Much more recent (and much less advanced) is the study of sequences of maps of the type $p_{n}\circ...\circ p_{1}$ where each $p_{n}$, $n\in\{1,2,...\},$ is a polynomial, but not necessarily the same one. Under the name of "generalized iteration", the works [5] and [6] deal with such sequences when the coefficients of underlying polynomials satisfy some conditions (see Subsection 2.3 for more details). For a recent study of the subject see [21]. In this article we extend the dynamical study to sequences of polynomials without necessarily assuming anything about their coefficients. Instead, we work with sequences of polynomials exhibiting certain asymptotic behavior. Our class includes not only the family $\mathcal{B}$ studied in [5] and [6], but also sequences of polynomials associated with a (fixed) compact polynomially convex set in the complex plane that have been widely studied in complex approximation and interpolation theory, often in connection with logarithmic potential theory ([27], [4], [9], [2]), namely sequences satisfying the Kalmár-Walsh condition. For examples see Subsection  2.3.
@@ 3. $(iii)$
We include a good amount of background material in our presentation to make it as self-contained as possible. Section 2 starts with preliminaries from logarithmic potential theory and complex approximation theory. Later, we introduce the notion of a guided sequence (such sequences satisfy conditions $(i)$ and $(ii)$ in the Main Theorem). Theorem 2.2.8 provides a characterization of guided sequences. Lemma 2.2.3 and Corollary 2.2.4 sharpen some previously available results. Examples of guided sequences appear in Subsection 2.3. Among them are Kalmár-Walsh sequences, important in approximation theory, and sequences from the class $\mathcal{B}$, previously studied in connection with generalized iteration. Using a concrete example, we point out that the class of guided sequences is larger than the class $\mathcal{B}$. More new material appears in Section 3, including the definition of a non-autonomous filled Julia set, as well as proofs of existence and properties of such sets. The approach using Klimek’s metric is explained. The presentation ends with Subsection 3.5, where we explore in more detail the case of the non-autonomous filled Julia set generated by $t_{d}=\frac{1}{2^{d-1}}T_{d},\ d\in\{1,2,...\}$, that is, by Chebyshev polynomials on the segment $[-1,1]$.
@@ 2.1. The Green function and related notions
Let us first define the Green function and the logarithmic capacity of a set (mostly following [25, Section II.4]). Our definitions are not the most general ones, but they still yield a notion of a regular set which agrees with the one considered in literature. In particular, finite sets are not regular.
@@
([26, 2.6]) Let $a\in\mathbb{C}$, $R>0$ and let $K=\overline{\mathbb{D}}(a,R):=\{z:|z-a|\leq R\}$. Then $g_{K}(z)=\log^{+}(|z-a|/R):=\max\{0,\log(|z-a|/R)\}$ and ${\rm cap}(K)=R$.
@@
([24, §1.2.b)]) $g_{[-1,1]}(z)=\log\left|z+\sqrt{z^{2}-1}\right|$, with the square root branch defined in $\mathbb{C}\setminus(-\infty,0)$ so that $\sqrt{1}=1$.
@@
([24, §1.2.c)]) Fix an $R>1$. Let $K=E_{R}$ be the ellipse with foci $-1,+1$ and semiaxes $a=\frac{1}{2}\bigl{(}R+\frac{1}{R}\bigr{)},\ b=\frac{1}{2}\bigl{(}R-\frac{1}{R}\bigr{)}$. Then $g_{K}(z)=\log^{+}\left({\left|z+\sqrt{z^{2}-1}\right|}/{R}\right)$ and ${\rm cap}(K)=(a+b)/2=R/2$.
@@ 3. $(iii)$
The Green function of a compact regular set $K\subset\mathbb{C}$ can be also characterized as follows (cf. [15, the beginning of Chapter 5 and Theorem 5.1.7]) for every $z\in\mathbb{C}$:
@@
If in addition $K$ is polynomially convex, the family $\{K_{\varepsilon}\}_{\varepsilon>0}$ forms a neighbourhood base of the set $K$ in $\mathbb{C}$ as shown in [16, Corollary 1]. It was proved by M. Mazurek (published in [26, Proposition 5.11]) that
@@
Let $K$ be a regular compact subset of $\mathbb{C}$. Then for every $\varepsilon>0$ the set $K_{\varepsilon}$ is polynomially convex and regular. Furthermore, ${\rm cap}(K_{\varepsilon})=\exp{(-\gamma+\varepsilon)}$.
@@ 2.2. Properties of sequences of polynomials
should hold “locally uniformly” in $\mathbb{C}\setminus{\rm conv}K$ (instead of in $\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)$). We prefer to write the explicit meaning of this notion, namely (**), following a personal communication with the authors.
@@ 2. $(ii)$
Observe that if $(p_{n})_{n=1}^{\infty}$ is a guided sequence, then for every mapping $\varphi:\mathbb{N}\longrightarrow\mathbb{N}$ such that $\varphi(1)=1$, the sequence $(p_{\varphi(n)})_{n=1}^{\infty}$ is also guided. If $\deg p_{1}>1$, the assumption $\varphi(1)=1$ can be omitted.
@@ 2.3. Examples of guided sequences of polynomials
When zeros of all $p_{n}$ do not have accumulation points in $\mathbb{C}\setminus K$, this is equivalent to $\lim\limits_{n\to\infty}|p_{n}(z)|^{1/n}={\rm cap}(K)\cdot\exp(g_{K}(z))$ uniformly on compact subsets of $\mathbb{C}\setminus K$ (cf. [4, Theorem 1.4], [27, Section 7.3, Theorem 3 and Section 7.4, Theorem 4] and [14, II.2.B, Theorem 1 and Lemma 1]). In particular, in this setting a KW sequence of polynomials associated with $K$ is $K$-guided, with $a=1$ and e.g. $b=-\log{\rm cap}(K)-1$. For the case when the zeros of $p_{n}$ are assumed to be uniformly bounded but are allowed to have limit points in the unbounded component of $\mathbb{C}\setminus K$ see [22] (cf. also the proof of Proposition 3.2.8).
@@
Now we turn to other examples. The authors of [5] define a class ${\mathcal{B}}$ of sequences $(f_{n})_{n=1}^{\infty}$, where $f_{n}(z)=\sum_{j=1}^{d_{n}}a_{n,j}z^{j}$ and $d_{n}\geq 2$ for any $n$. The coefficients of polynomials in sequences in $\mathcal{B}$ satisfy some conditions, in particular
@@
It was shown in [5, 4.1] that the condition $(iii)$ in our Theorem 2.2.8 holds for sequences of class $\mathcal{B}$. Therefore they are guided. However, our methods apply also to some sequences not belonging to $\mathcal{B}$, because the condition (P2) does not hold for certain KW sequences of polynomials.
@@
The sequence of the Chebyshev polynomials on $[-1,1]$ does not belong to the class $\mathcal{B}$.
@@
Thus for the polynomial $\frac{1}{2^{n-1}}T_{n}$ the coefficient corresponding to $z^{n-2}$ is $-n/4$. Hence condition (P2) does not hold and $\left(\frac{1}{2^{n-1}}T_{n}\right)_{n=1}^{\infty}\notin\mathcal{B}.$
For the explicit expression of coefficients of $T_{n}$ corresponding to the powers of the variable $z$ see e.g. [23, T1.2 (40)].
@@
Let $L\subset\mathbb{C}$ be any compact regular polynomially convex set such that $K\cap L=\emptyset$ and $\mathbb{C}\setminus(K\cup L)$ is connected. Take $\delta>0$ such that $K\cup L\subset\overline{\mathbb{D}}(0,\delta)$. Let $(p_{n})_{n=1}^{\infty}$ and $(q_{n})_{n=1}^{\infty}$ be KW sequences of polynomials associated with $K$ and $L$ respectively, such that all zeros of $p_{n}$ lie in $K$ and all zeros of $q_{n}$ lie in $L$. Let further $k_{n},l_{n}$ be two sequences of integers such that $0<k_{n}<l_{n}$, $n\in\{1,2,...\}$, and $a=\lim_{n_{\to}\infty}k_{n}/l_{n}$. If $a=k/l\in\mathbb{Q}$ with $k,l$ coprime, simply take the sequences $k_{n},l_{n}$ to be constant, $k_{n}=k$, $l_{n}=l$. If $a\notin\mathbb{Q}$, take $l_{n}\longrightarrow\infty$ as $n\to\infty$ (e.g., by fixing a decimal expansion of $a$).
@@ 3.1. Autonomous Julia sets
Let us define the notion of escape radius of a polynomial, following [19, page 53] (a slightly different definition was proposed in [10]).
@@ 3.2. Non-autonomous Julia sets
We will consider now a generalization of the autonomous Julia set. As in Definition 3.1.1, here also usually only polynomials of degree greater than 1 are considered (cf. e.g. [6], [19]). However, since we are interested in particular in KW sequences of polynomials, we will allow one polynomial to be of degree  1.
@@
If $(p_{n})_{n=1}^{\infty}$ in Definition 3.2.1 above is periodic (i.e., there exists an $m$ such that $p_{m+i}=p_{i}$ for every $i$), then we obtain the autonomous filled Julia set (see Definition 3.1.1) of the polynomial $p_{m}\circ....\circ p_{1}$. In particular for a constant sequence we also get an autonomous Julia set. The interesting case is when the sequence is not periodic, e.g., when each polynomial in the sequence has a different degree.
@@
Recall that the Julia set obtained in this theorem may be finite, as shown in Example 3.2.3. Our aim is to find regular sets, therefore we need an additional assumption. Remark 2.3.5 shows that the approach from [5] may be insufficient in our study. Another course of action was proposed in [19], where it was required that all polynomials from the sequence $(p_{n})_{n=1}^{\infty}$ have a common escape radius $R$ such that
@@ 3.3. Klimek’s metric
$\forall n\in\mathcal{N}\ \forall\zeta\in\partial\overline{\mathbb{D}}(0,C_{3}^{\deg p_{n}}):\quad\frac{1}{\deg p_{n}}g_{E}(\zeta)\leq g_{E}(\zeta)\leq C_{1}.$
@@
Combining this with (3.9) gives
(3.11) $\forall n\notin\mathcal{N}:\quad\sup_{z\in E}g_{p_{n}^{-1}(E)}(z)\leq C_{4}:=C_{2}-\log{\rm cap}(E)+\varepsilon.$

### v3

@@ 1. Introduction
Behavior of iterates $p^{n}:=p^{\circ n}$ of a polynomial $p:\mathbb{C}\longrightarrow\mathbb{C}$ has been intensely studied since the nineteenth century. For an introduction to the subject see e.g. [7]. Much more recent (and much less advanced) is the study of sequences of maps of the type $p_{n}\circ...\circ p_{1}$ where each $p_{n}$, $n\in\{1,2,...\},$ is a polynomial, but not necessarily the same one. Under the name of "generalized iteration", the works [5] and [6] deal with such sequences when the coefficients of underlying polynomials satisfy some conditions (see Subsection 2.3 for more details). For a recent study of iteration of quadratic polynomials with random coefficients see [21] and [22].
In this article we extend the dynamical study to sequences of polynomials without necessarily assuming anything about their coefficients. Instead, we work with sequences of polynomials exhibiting certain asymptotic behavior, called guided sequences. Our class includes not only the family $\mathcal{B}$ studied in [5] and [6], but also important sequences of polynomials associated with a (fixed) compact polynomially convex set in the complex plane, namely sequences satisfying the Kalmár-Walsh condition. Such sequences have been widely studied in complex approximation and interpolation theory, often in connection with logarithmic potential theory ([29], [4], [9], [2]). For examples see Subsection  2.3.
@@ 3. $(iii)$
We include a considerable amount of background material in our presentation to make it as self-contained as possible. Section 2 starts with preliminaries from logarithmic potential theory and complex approximation theory. Later, we introduce the notion of a guided sequence (such sequences satisfy conditions $(i)$ and $(ii)$ in the Main Theorem). Theorem 2.2.8 provides a characterization of guided sequences. Lemma 2.2.3 and Corollary 2.2.4 sharpen some previously available results. Examples of guided sequences appear in Subsection 2.3. Among them are Kalmár-Walsh sequences, important in approximation theory, and sequences from the class $\mathcal{B}$, previously studied in connection with generalized iteration. Using concrete examples (see Remark 2.3.5 and Example 2.3.6) we discuss relations between these classes and the class of guided sequences. More new material appears in Section 3, including the definition of a non-autonomous filled Julia set, as well as proofs of existence and properties of such sets. The approach using Klimek’s metric is explained. The presentation ends with Subsection 3.5, where we explore in more detail the case of the non-autonomous filled Julia set generated by $t_{d}=\frac{1}{2^{d-1}}T_{d},\ d\in\{1,2,...\}$, that is, by Chebyshev (minimal) polynomials on the segment $[-1,1]$.
@@ 2.1. The Green function and related notions
Let us first define the Green function and the logarithmic capacity of a set (mostly following [26, Section II.4]). Our definitions are not the most general ones, but they still yield a notion of a regular set which agrees with the one considered in literature. In particular, finite sets are not regular.
@@
([27, 2.6]) Let $a\in\mathbb{C}$, $R>0$ and let $K=\overline{\mathbb{D}}(a,R):=\{z:|z-a|\leq R\}$. Then $g_{K}(z)=\log^{+}(|z-a|/R):=\max\{0,\log(|z-a|/R)\}$ and ${\rm cap}(K)=R$.
@@
([25, §1.2.b)]) $g_{[-1,1]}(z)=\log\left|z+\sqrt{z^{2}-1}\right|$, with the square root branch defined in $\mathbb{C}\setminus(-\infty,0)$ so that $\sqrt{1}=1$.
@@
([25, §1.2.c)]) Fix an $R>1$. Let $K=E_{R}$ be the ellipse with foci $-1,+1$ and semiaxes $a=\frac{1}{2}\bigl{(}R+\frac{1}{R}\bigr{)},\ b=\frac{1}{2}\bigl{(}R-\frac{1}{R}\bigr{)}$. Then $g_{K}(z)=\log^{+}\left({\left|z+\sqrt{z^{2}-1}\right|}/{R}\right)$ and ${\rm cap}(K)=(a+b)/2=R/2$.
@@ 3. $(iii)$
The Green function of a compact regular set $K\subset\mathbb{C}$ can be also characterized (cf. [15, the beginning of Chapter 5 and Theorem 5.1.7]) for every $z\in\mathbb{C}$ as follows:
@@
If in addition $K$ is polynomially convex, the family $\{K_{\varepsilon}\}_{\varepsilon>0}$ forms a neighbourhood base of the set $K$ in $\mathbb{C}$ as shown in [16, Corollary 1]. It was proved by M. Mazurek (published in [27, Proposition 5.11]) that
@@
Let $K$ be a regular compact subset of $\mathbb{C}$. Then for every $\varepsilon>0$ the set $K_{\varepsilon}$ is polynomially convex and regular. Furthermore, ${\rm cap}(K_{\varepsilon})=\exp{(-\gamma+\varepsilon)}=\exp(\varepsilon){\rm cap}(K)$.
@@ 2.2. Properties of sequences of polynomials
should hold “locally uniformly” in $\mathbb{C}\setminus{\rm conv}K$ (instead of in $\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)$). We prefer to write the explicit meaning of this notion, namely (**), following personal communication with the authors.
@@ 2. $(ii)$
Observe that if $(p_{n})_{n=1}^{\infty}$ is a guided sequence, then for every mapping $\varphi:\mathbb{N}\longrightarrow\mathbb{N}$ such that $\varphi(k)>1$ for all $k>1$ , the sequence $(p_{\varphi(n)})_{n=1}^{\infty}$ is also guided. If $\deg p_{1}>1$, the assumption $\varphi(k)>1$ for all $k>1$ can be omitted.
@@ 2.3. Examples of guided sequences of polynomials
When zeros of all $p_{n}$ do not have accumulation points in $\mathbb{C}\setminus K$, this is equivalent to $\lim\limits_{n\to\infty}|p_{n}(z)|^{1/n}={\rm cap}(K)\cdot\exp(g_{K}(z))$ uniformly on compact subsets of $\mathbb{C}\setminus K$ (cf. [4, Theorem 1.4], [29, Section 7.3, Theorem 3 and Section 7.4, Theorem 4] and [14, II.2.B, Theorem 1 and Lemma 1]). In particular, in this setting a KW sequence of polynomials associated with $K$ is $K$-guided, with $a=1$ and e.g. $b=-\log{\rm cap}(K)-1$. For the case when the zeros of $p_{n}$ are assumed to be uniformly bounded but are allowed to have limit points in the unbounded component of $\mathbb{C}\setminus K$ see [23] (cf. also the proof of Proposition 3.2.8).
@@
Now we turn to other examples. The authors of [5] introduced a class $\mathcal{B}$ of sequences $(f_{n})_{n=1}^{\infty}$, where $f_{n}(z)=\sum_{j=1}^{d_{n}}a_{n,j}z^{j}$ and $d_{n}\geq 2$ for any $n$. Formally, in [5, 4.1] $\mathcal{B}$ was defined as the class of sequences of polynomials satisfying a variant of the condition $(iii)$ in our Theorem 2.2.8. Therefore such sequences are guided. However, in practice, in [5] as well as in other works building on it, only sequences of polynomials satisfying some sufficient condition were considered to be members of $\mathcal{B}$. In particular, they would satisfy
@@
The necessity of condition (P2) for membership in $\mathcal{B}$ was not addressed in the existing literature. We would like to point out that (P2) does not hold for certain KW sequences of polynomials. Therefore the class of guided sequences is larger than the class of sequences satisfying (P2). Here is an explicit example.
@@
The sequence of the Chebyshev polynomials on $[-1,1]$ does not satisfy (P2).
@@
Thus for the polynomial $\frac{1}{2^{n-1}}T_{n}$ the coefficient corresponding to $z^{n-2}$ is $-n/4$. Hence condition (P2) does not hold. ∎
For the explicit expression of coefficients of $T_{n}$ corresponding to the powers of the variable $z$ see e.g. [24, T1.2 (40)].
@@
Let $L\subset\mathbb{C}$ be any compact regular polynomially convex set such that $K\cap L=\emptyset$ and $\mathbb{C}\setminus(K\cup L)$ is connected. Take $\delta>0$ such that $K\cup L\subset\overline{\mathbb{D}}(0,\delta)$. Let $(p_{n})_{n=1}^{\infty}$ and $(q_{n})_{n=1}^{\infty}$ be KW sequences of polynomials associated with $K$ and $L$ respectively, such that all zeros of $p_{n}$ lie in $K$ and all zeros of $q_{n}$ lie in $L$. Let further $k_{n},l_{n}$ be two sequences of integers such that $0<k_{n}<l_{n}$, $n\in\{1,2,...\}$, and $a=\lim_{n_{\to}\infty}k_{n}/l_{n}$. If $a=k/l\in\mathbb{Q}$ with $k,l$ coprime, simply take the sequences $k_{n},l_{n}$ to be constant, $k_{n}=k$, $l_{n}=l$. If $a\notin\mathbb{Q}$, take $l_{n}\longrightarrow\infty$ as $n\to\infty$ (e.g., fixing an approximation of $a$ by decimal fractions).
@@ 3.1. Autonomous Julia sets
Let us define the notion of an escape radius of a polynomial, following [19, page 53] (a slightly different definition was proposed in [10]).
@@ 3.2. Non-autonomous Julia sets
We will consider now a generalization of the autonomous Julia set. Even though usually only polynomials of degree greater than 1 are considered (cf. e.g. [6], [19]), we will allow one polynomial to be of degree 1. This in particular takes into account KW sequences of polynomials and enables their study.
@@
If $(p_{n})_{n=1}^{\infty}$ in Definition 3.2.1 above is periodic (i.e., there exists an $m$ such that $p_{m+i}=p_{i}$ for every $i$), then we obtain the autonomous filled Julia set (see Definition 3.1.1) of the polynomial $p_{m}\circ....\circ p_{1}$. In particular for a constant sequence we also get the autonomous Julia set. The interesting case is when the sequence is not periodic, e.g., when each polynomial in the sequence has a different degree.
@@
Recall that the Julia set obtained in this theorem may be finite, as shown in Example 3.2.3. Our aim is to find regular sets, therefore we need an additional assumption. Remark 2.3.5 shows that the approach from [5], where bounds on coefficients were assumed, may be insufficient in our study. Another course of action was proposed in [19], where it was required that all polynomials from the sequence $(p_{n})_{n=1}^{\infty}$ have a common escape radius $R$ such that
@@ 3.3. Klimek’s metric
$\forall n\in\mathcal{N}\ \forall\zeta\in\partial\overline{\mathbb{D}}\left(0,C_{3}^{\deg p_{n}}\right):\quad\frac{1}{\deg p_{n}}g_{E}(\zeta)\leq g_{E}(\zeta)\leq C_{1}.$
@@
Recall that $\log C_{3}=C_{2}>0$ and $\varepsilon-\log{\rm cap}(E)>0$ by the choice of $\varepsilon$. Combining this with (3.9) gives
$\forall n\notin\mathcal{N}:\quad\sup_{z\in E}g_{p_{n}^{-1}(E)}(z)$ $\leq C_{2}+\frac{1}{\deg p_{n}}(\varepsilon-\log{\rm cap}(E))$
(3.11) $\leq C_{2}+\varepsilon-\log{\rm cap}(E)=:C_{4}.$
@@ 3.4. Julia sets for guided sequences
Let us note in passing that if $f_{n}:\mathbb{C}\ni z\longmapsto z^{2}+c_{n}\in\mathbb{C}$ for $n\geq 1$ with $(c_{n})_{n=1}^{\infty}\subset\overline{\mathbb{D}}(0,\frac{1}{4})$, then $(f_{n})_{n=1}^{\infty}$ is guided and $g_{{\mathcal{K}}[(f_{n})_{n=1}^{\infty}]}$ is Hölder continuous (cf. [28]).